OpenStudy (anonymous):

A rubber ball dropped on a hard surface takes a sequence of bounces, each one 4/5 as high as the preceding one. If this ball is dropped from a height of 10 feet, what is the total vertical distance it has traveled after it hits the surface the fifth time? Be sure to account for both drop and rebound distances. We are learning sequences and series I'm confused on how to do this >.<

OpenStudy (anonymous):

I just learned this in my class, I am pretty sure I can help you, let me go find my notes on it.

OpenStudy (anonymous):

Ok :)

OpenStudy (shiraz14):

This is a hybrid G.P. with T1=8, r=0.8, n=4, (excluding the first fall of 10 feet). i.e. to say: Total vertical distance = [2x(Sum of G.P. where T1=8, r=0.8, n=4)]+10 feet.

OpenStudy (anonymous):

ok so if the height is 10 and the rebound is 4/5 we want to first find the ratio between the sequences

OpenStudy (anonymous):

\[10(ratio)^{i-1}\sum_{i-1}^{5} \frac{ 4 }{ 5 }(ratio)^{i-1}\]would be the formula set up we would want to use after we found the ratio

OpenStudy (anonymous):

Wouldn't the ratio be 4/5 ? "4/5 as high as the preceding one"

OpenStudy (anonymous):

oh you are right, so 4/5 would go in the part where I have ratio, and 4/5 of 10 would go were I have 4/5

OpenStudy (anonymous):

since we are not going to infinity and just five it would be i not i-1. so the rebound would be 4/5 or 8/10 the first bounce, so the equation would look like\[10(\frac{ 4 }{ 5 })^{i}\sum_{i}^{5}\frac{ 4 }{ 5}(\frac{ 4 }{ 5})^{i}\]

OpenStudy (anonymous):

\[10(\frac{ 1 }{ 1-\frac{ 4 }{ 5 } })-5(\frac{ 1 }{ 1-\frac{ 4 }{ 5 } })\]

OpenStudy (anonymous):

50-15=45 ft

OpenStudy (anonymous):

does that look right?

OpenStudy (anonymous):

Not really? I'm confused to be honest >.<

OpenStudy (anonymous):

k what do you understand, and I'll pick up from there

OpenStudy (anonymous):

Hold on a second let me reread your stuff :)

OpenStudy (anonymous):

Okay doesn't \[\sum_{n=1}^5 10 (\frac{4}{5})^n \] that work?

OpenStudy (anonymous):

*why doesn't that work?

OpenStudy (anonymous):

That can, I used it so that we had a1 and a2, but if we wanted to work with just a1 that is probably a simpler approach. I think I worked it out to how far the ball traveled by the time it came to a stop anyways, and not the fifth bounce. Looking back over my notes I think it would be easier to use the formula\[S _{n}=a _{1}\frac{ 1-r ^{n} }{ 1-r }\] where S=total distance traveled r=10 and n=5

OpenStudy (anonymous):

well a= 10 and r= 4/5

OpenStudy (anonymous):

That gives me: 33.616 + 33.616 = 67.232 Which isn't one of my answer choices :/

OpenStudy (anonymous):

\[S _{5}=10\frac{ 1-\frac{ 4 }{ 5 }^{5} }{ 1-\frac{ 4 }{ 5 } }\]

OpenStudy (anonymous):

what are your answer choices?

OpenStudy (anonymous):

A 28 77/125 B 49 1/25 C 57 29/125 D 59 29/125

OpenStudy (anonymous):

well we could always make a table, and go about it without an equation

OpenStudy (anonymous):

Uh not sure what I exactly would do with a table :/ explain please?

OpenStudy (anonymous):

well each time the ball drops it comes back up 4/5 distance, so it drops ten, comes up 4/5 which means it didn't not come back up 1/5 of the distance so 10-1/5 is 49/5. The table shows each distance of fall, with each distance of rebound. Then we add them all together to get the answer.

OpenStudy (anonymous):

I think the measurements on the table are a little off. Let me try to convert them to decimals, see if I got it wrong some where.

OpenStudy (anonymous):

nope i keep getting 86, so I'm not sure. THe formula Shiraz14 gave doesn't make sense to me either. I do know... I thought I could figure it out, but obviously I need more practice with these problems. Sorry :(